OpenStudy (anonymous):

A school system is reducing the amount of dumpster loads of trash removed each week. In week 5, there were 40 dumpster loads of waste removed. In week 10, there were 30 dumpster loads removed. Assume that the reduction in the amount of waste each week is linear. Write an equation in function form to show the amount of trash removed each week. f(x) = - 2x + 40 f(x) = 2x + 40 f(x) = - 2x + 50 f(x) = 2x + 50

OpenStudy (aric200):

y2-y1/x2-x1 30-40/10-5 -10/5 -2 y=-2x+b 30=-2*10+b 30=-20+b 50=b y=-2x+50

OpenStudy (aric200):

C.